OpenStudy (anonymous):

How do you solve 2sinx-1=0 without a calculator?

OpenStudy (anonymous):

\[2sinx = 1\] \[sin(x) = \frac{1}{2}\] What value of sin is 1/2??

OpenStudy (anonymous):

sinx=1/2 so sinx=sin30 so x=30 degrees

OpenStudy (diyadiya):

Add 1 on both sides 2sinx = 1 divide 2 on both sides sinx = 1/2 x= \(sin^{-1} \frac{1}{2} \)

OpenStudy (anonymous):

do you have to draw a graph of sin to find these values for sinx=0.5?

OpenStudy (anonymous):

Thank you for such quick replies :)

OpenStudy (anonymous):

So, \[sin(x) = sin ( \frac { \pi }{6}) \implies x = \frac{\pi}{6}\]

OpenStudy (cwrw238):

you can also commit to memory this triangle